40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 21 16:51:43 2018, gautam, Update, VAC, c1vac1 is unresponsive 
    Reply  Mon Sep 24 15:06:10 2018, gautam, Update, VAC, c1vac1 reboot + TP1 controller replacement beforeReboot.pngafterReboot.pngCC1.png
       Reply  Wed Sep 26 10:07:16 2018, Steve, Update, VAC, why reboot c1vac1 NOcomm.pngReboot_&_sawp.pngc1vac1&2_.jpg
Message ID: 14217     Entry time: Wed Sep 26 10:07:16 2018     In reply to: 14215
Author: Steve 
Type: Update 
Category: VAC 
Subject: why reboot c1vac1 

Precondition: c1vac1 & c1vac2 all LED warning lights green [ atm3 ], the only error message is in the gauge readings NO COMM, dataviewer will plot zero [ atm1 ], valves are operational

When our vacuum gauges read " NO COMM " than our INTERLOCKS  do  NOT communicate either.

So V1 gate valve and PSL output shutter can not be triggered to close if the the IFO pressure goes up.                        

   [ only CC1_HORNET_PRESSURE reading is working in this condition because it goes to a different compuer ] 

Quote:

[steve, gautam]

Following the procedure in this elog, we effected a reset of the vacuum slow machines. Usually, I just turn the key on these crates to do a power cycle, but Steve pointed out that for the vacuum machines, we should only push the "reset" button.

While TP1 was spun down, we took the opportunity to replace the TP1 controller with a spare unit the company has sent us for use while our unit is sent to them for maintenance. The procedure was in principle simple (I only list the additional ones, for the various valve closures, see the slow machine reset procedure elog):

  • Turn power off using switch on rear.
  • Remove 4 connecting cables on the back.
  • Switch controllers.
  • Reconnect 4 cables on the back panel.
  • Turn power back on using switch on rear.

However, we were foiled by a Philips screw on the DB37 connector labelled "MAG BRG", which had all its head worn out. We had to make a cut in this screw using a saw blade, and use a "-" screwdriver to get this troublesome screw out. Steve suspects this is a metric gauge screw, and will request the company to send us a new one, we will replace it when re-installing the maintaiend controller. 

Attachments #1 and #2 show the Vacuum MEDM screen before and after the reboot respectively - evidently, the fields that were reading "NO COMM" now read numbers. Attachment #3 shows the main volume pressure during this work.

Quote:

The problem will be revisited on Monday.

 

Attachment 1: NOcomm.png  23 kB  Uploaded Wed Sep 26 11:10:10 2018  | Show | Hide all | Show all
Attachment 2: Reboot_&_sawp.png  51 kB  Uploaded Wed Sep 26 11:10:52 2018  | Hide | Hide all | Show all
Reboot_&_sawp.png
Attachment 3: c1vac1&2_.jpg  297 kB  Uploaded Thu Sep 27 09:29:05 2018  | Show | Hide all | Show all
ELOG V3.1.3-